# GPU Acceleration of SQL Analytics on Compressed Data

### v1

@@ 5. Logical Operations
Table 3. Output MaskColumn encoding for AND Operation. The table is symmetric and we only show the upper triangular matrix for better readability.
@@ 5.1. AND
  • Plain mask AND Plain mask: This is a straightforward boolean operation using PyTorch’s logical AND operator (&). For example, if we have two Plain masks $[T,F,T,F]$ and $[T,T,F,F]$, the result would be $[T,F,F,F]$.
@@
  • RLE mask AND Plain mask: Directly performing AND between an RLE mask and a Plain mask is inefficient because it requires searching for each ‘True’ value in the Plain mask. A more efficient approach is to convert the RLE mask into either an Index mask (rle_to_index) or a Plain mask (rle_to_plain), and then apply the AND operation following Table 2. The choice depends on the sparsity of the RLE mask: if sparse, convert to an Index mask; otherwise, convert to a Plain mask.
  • RLE mask AND Index mask: This operation checks whether each index value in the Index mask falls within any range specified by the RLE mask. PyTorch’s bucketize function provides native support for this step. Depending on the relative sizes of the masks, we either use the Index mask to bucketize the RLE mask (idx_in_rle) or use the RLE mask to bucketize the Index mask (rle_contain_idx).
  • Plain mask AND Index mask: This operation selects index values where the corresponding positions in the Plain tensor are ‘True’, using PyTorch’s subscript operator ([][][ ]).
  • Index mask AND Index mask: This operation checks whether each value in one index tensor also appears in the other one. This can be efficiently performed using bucketization, with the larger tensor chosen as the bucketized input.
@@
$\left\{\begin{array}[]{c}\text{{merge\_sorted\_idx}($m_{1}$, {\color[rgb]{0.3,0.3,1}$m_{2}$}) }\\ \text{{merge\_sorted\_idx}({\color[rgb]{0.3,0.3,1}$m_{2}$}, $m_{1}$) }\\ \text{{concat\_sort}($m_{1}$, {\color[rgb]{0.3,0.3,1}$m_{2}$}) }\end{array}\right.$
@@
  • Plain mask OR Plain mask: This is a simple boolean OR operation using PyTorch’s logical OR operator (||||). For example, if we have two Plain masks $[T,F,T,F]$ and $[T,T,F,F]$, the result would be $[T,T,T,F]$.
  • RLE mask OR RLE mask: This operation computes the union of two sorted lists of ranges. We identify the start and end points of consecutive True values by computing differences. Non-overlapping segments are then determined using a mask.
  • RLE mask OR Plain mask: To perform this operation efficiently, we first convert the RLE mask into either an Index mask (rle_to_index) or a Plain mask (rle_to_plain), and then apply the OR operation. The choice depends on the sparsity of the RLE mask: if sparse, convert to an Index mask; otherwise, convert to a Plain mask.
  • RLE Mask OR Index Mask: This involves bucketizing based on the relative sizes of the RLE and Index masks. We use either the Index mask to bucketize the RLE mask (idx_in_rle) or use the RLE mask to bucketize the Index mask (rle_contain_idx).
  • Plain Mask OR Index Mask: This operation is performed using PyTorch’s subscript ([]) operator.
  • Index Mask OR Index Mask: This operation merges two sorted Index tensors (merge_sorted_idx). For efficiency, we bucketize the larger tensor. A flag tracks the origin of each element, and the final merge is performed using a conditional selection (‘where’). If both Index tensors are small, the most efficient approach is to simply concatenate and sort them.
@@ 5.3. NOT
Finally, we describe the NOT logical operator, which takes a single MaskColumn as input and produces an output MaskColumn.
  • NOT Plain mask: This is a straightforward bitwise negation using PyTorch’s complement ($\sim$) operator.
  • NOT RLE mask: We use the complement_rle primitive. The key idea is to compute the gaps between consecutive runs. Each gap starts at the end of the previous run plus one and ends at the start of the next run minus one. This requires an additional metadata to track the total number of the rows in the column.
  • NOT Index mask: We use the complement_index primitive. Similar to the RLE case, we compute gaps between indices, and need to keep track of the column size. Notably, the output is in RLE format rather than an Index representation. This is because Index MaskColumns are assumed to be sparse, making their NOT result more continuous and thus well-suited for RLE.
Figure 3. Illustration of NOT logical operator.
Example 5.
  • NOT on RLE mask. For an RLE mask with $s=[0,4]$, $e=[1,6]$, and $total\_size=8$, we apply the complement_rle primitive. First, we compute the potential start points, $s^{\prime}$, for inverted ranges by concatenating $[-1]$ with the original end values and adding 1. $s^{\prime}=[0,2,7]$. Then, we compute potential end points, $e^{\prime}$, by concatenating the original start values with $[total\_size]$ and subtracting 1. $e^{\prime}=[-1,3,7]$. We keep only the valid ranges where $s^{\prime}\leq e^{\prime}$, which gives us $s^{\prime}=[2,7]$ and $e^{\prime}=[3,7]$. This is an RLE mask covering the gaps between the original runs.
  • Step 3: NOT on Index mask. For an Index mask with $p=[2,5]$ and $total\_size=8$, we apply the complement_index primitive. We compute potential start points by concatenating $[-1]$ with the original indices and adding 1. $s=[0,3,6]$. We compute potential end points by concatenating the original indices with $[total\_size]$ and subtracting 1. $e=[1,4,7]$. After applying a validity mask to ensure $s\leq e$, $s<total\_size$, and $e\geq 0$, all position ranges remain valid. This is an RLE Mask with runs covering all positions not in the original Index mask.
@@ 5.4. Operating on Composite MaskColumns
To apply the previously described logical operators to Composite (Plain + Index, or RLE + Index) masks, we extend our approach accordingly. A key observation is that a Composite MaskColumn layout is essentially an OR operation between mask tensors. This allows us to leverage Boolean algebra (De Morgan’s Law, Associative and Distributive rules) to rewrite logical operations on Composite masks. Let $m(RLE)$ and $m(Index)$ represent the RLE and Index tensors of Composite mask $m$.
  • NOT: $\neg(m_{1}(RLE)\vee m_{1}(Index))=(\neg m_{1}(RLE))\wedge(\neg m_{1}(Index))$
  • OR: $({\color[rgb]{0.3,0.3,1}m_{1}(RLE)}\vee{\color[rgb]{0.3,0.3,1}m_{1}(Index)})\vee({\color[rgb]{1,0.3,0.3}m_{2}(RLE)}\vee{\color[rgb]{1,0.3,0.3}m_{2}(Index)})$
    $=({\color[rgb]{0.3,0.3,1}m_{1}(RLE)}\vee{\color[rgb]{1,0.3,0.3}m_{2}(RLE)})\vee({\color[rgb]{0.3,0.3,1}m_{1}(Index)}\vee{\color[rgb]{1,0.3,0.3}m_{2}(Index)})$
  • AND:$({\color[rgb]{0.3,0.3,1}m_{1}(RLE)}\vee{\color[rgb]{0.3,0.3,1}m_{1}(Index)})\wedge({\color[rgb]{1,0.3,0.3}m_{2}(RLE)}\vee{\color[rgb]{1,0.3,0.3}m_{2}(Index)})$
    $=({\color[rgb]{0.3,0.3,1}m_{1}(RLE)}\wedge{\color[rgb]{1,0.3,0.3}m_{2}(RLE)})\vee({\color[rgb]{0.3,0.3,1}m_{1}(RLE)}\wedge{\color[rgb]{1,0.3,0.3}m_{2}(Index)})\vee({\color[rgb]{0.3,0.3,1}m_{1}(Index)}\wedge{\color[rgb]{1,0.3,0.3}m_{2}(RLE)})\vee({\color[rgb]{0.3,0.3,1}m_{1}(Index)}\wedge{\color[rgb]{1,0.3,0.3}m_{2}(Index)})$
@@ 6. Arithmetic, Comparison, and Selection Operations
Next, we focus on a general class of operators whose results are a point-wise/position-wise operation on values from two input columns, for positions that are common to both columns. This includes the logical AND operator that we discussed in Section 5.1, binary arithmetic operators, comparison operators, and selection operators using a filter predicate. Joins need additional handing that we discuss in Section 8. The logical OR operator that we discussed in Section 5.2 does not fit into this category in the general case since the result covers a union of the set of positions from the input columns which can be larger than the set of positions for any one column in the presence of gaps (e.g., due to a prior filter operation being applied to the column, and the result not being compacted thereafter).
The main challenge in performing these operations on compressed inputs is that, in contrast to Plain encoding, there is no longer an alignment between positional representations of the inputs. Thus, point-wise operations are not directly possible. Instead, we need to first align the positional representations of the inputs, including the value tensors as needed, and then perform the operations on the aligned segments. We call this an Alignment operation.
@@
Consider the operation $A+B$ on input columns $A$ and $B$ which are both in RLE format. $A.v=[4,1,3],A.s=[0,10,20],A.e=[9,19,39]$; $B.v=[6,8],B.s=[0,15],B.e=[14,39]$. Thus, $A$ has 3 runs of lengths 10, 10, and 20 and $B$ has 2 runs of lengths 15 and 25. Similar examples apply for other operators such as ,,/,<,>,=-,*,/,<,>,=- , ∗ , / , < , > , =, etc.
Note that due to the misalignment of runs, we cannot directly do $A.v+B.v$. Although in this example the two columns have different runs, misalignment can happen even with the same number of runs but differences in run lengths for at least one run.
To perform this operation, we first align the runs of the input columns. This is exactly how we handle the AND operator (Section 5.1) with an additional step of reconstructing the value ($v$) tensors for the result. In this example, the result of the alignment operation would be two RLE columns $R_{1}$ and $R_{2}$ with $R_{1}.s=R_{2}.s=[0,10,15,20]$, $R_{1}.e=R_{2}.e=[9,14,19,39]$, $R_{1}.v=[4,1,1,3]$, $R_{2}.v=[6,6,8,8]$. Both columns have identical $s$ and $e$ tensors. We can now do a point-wise addition of the value tensors to get the value for the 4 runs in the result: $R_{1}.v+R_{2}.v=[10,7,9,11]$.
The operation is simple if one of the operands is a scalar expression or literal, e.g., $A*2,1-A,A\geq 3$, etc. in which case we do not need to align anything but instead just operate on the value tensors for both RLE and Index encoding.
For selection operations with a filter predicate, e.g., SELECT A FROM D WHERE P, we first compute a MaskColumn $m$ according to the evaluation of predicate P, then align $m$ and A, and finally apply $m$ to the tensor representations of $A$ (if needed). Since $m$ only tracks True values for RLE and Index encoding, the alignment step itself causes the selection and the final application step is not needed for those cases. However, if $m$ or $A$ is Plain encoded, then the final mask application step is needed to get the result.
@@ 7. Group-By-Aggregation Operations
For aggregation queries, the process is decomposed into two phases.
@@
Both phases are conceptually straightforward in PyTorch—function torch.unique can handle grouping by building inverse indices based on the unique values in the group-by columns, and function torch.scatter can handle aggregation based on these inverse indices. However, the challenge lies in managing heterogeneous compression schemes across DataColumns. For example, consider a query that groups by two columns, one of which is RLE-compressed and the other is Plain. The torch.unique function will not work because it requires two Plain DataColumns, but the values of the RLE column are not aligned with those of the Plain column. This heterogeneity prevents the direct application of torch.unique and torch.scatter due to misalignment in the data representations. To solve this problem, we apply our novel Alignment technique, described in Section 6, to both the Grouping and Aggregating phases.
@@ 7.1. Grouping Phase
The grouping phase constructs an inverse index that maps each row to its corresponding group; rows in the same group share the same data values across the group-by columns. It takes a set of aligned DataColumns for the group-by (potentially non-compact) as input and returns an inverse index. The inverse index is a single tensor from 0 to N-1, where N is the number of unique values across the group-by columns. The length of the inverse index is the same as the number of values in the group-by columns—for Plain data, it is the number of rows; for RLE and Index data, it is the number of RLE runs or index points.
@@ 7.2. Aggregating Phase
To implement aggregation, we apply the torch.scatter function to the data columns, scattering based on the inverse indices. This approach efficiently computes aggregates like SUM, COUNT, AVG, STD, VAR, MIN, and MAX for each group. Note, however, that if the data is RLE-compressed, each value in the data column must be repeated based on its corresponding RLE range size. Therefore, we need to account for this when scattering the data. This depends on the aggregation function.
@@
  • SUM, COUNT: For RLE columns, instead of expanding the data and counting/summing individual values, we compute the run lengths: $l=e-s+1$. This is sufficient for COUNT. For SUM, we multiply the run lengths by the values ($v*l$).
@@
Example 7.
Consider a SQL query ”SELECT SUM(B) GROUP BY A” on RLE-compressed data as shown in Figure 4. The input consists of a group-by column A with values $[A,B,A]$ spanning runs $[0-1]$, $[2-4]$, $[5-8]$ and a sum column B with values $[0,3,8]$ spanning the same runs $[0-1]$, $[2-4]$, $[5-8]$.
  • Step 1: Grouping. We apply the grouping operation on the group-by column to identify unique groups. Using torch.unique on the values $[A,B,A]$, we identify two unique groups (A and B) and generate an inverse index, $x=[0,1,0]$ that maps each run to its corresponding group: the first run (group-by value A) maps to group 0, the second run (group-by value B) maps to group 1, and the third run (group-by value A) again maps to group 0. This creates a temporary RLE column $g$ with $v=x$ and with $s$ and $e$ the same as that of the group-by column ($A$).
  • Step 2: Alignment. Next, we align $g$ and the aggregation column ($B$). After alignment, as shown in the figure, both columns share the same runs $[0-1]$, $[2-4]$, and $[5-8]$. The aggregation column B is represented with aligned values $[3,3,3]$ for these runs.
  • Step 3: Aggregation. We do this in two steps—(a) perform a parallel aggregation for each segment in the aligned aggregation column by computing $v\times l$ using its values $[3,3,3]$ and run lengths $[2,3,4]$, resulting in $[3\times 2,3\times 3,3\times 4]=[6,9,12]$, (b) perform a parallel aggregation for each group-by key by applying torch.scatter on the tensor $[6,9,12]$ using the inverse index $[0,1,0]$. The scatter results in values $6+12=18$ (for group A, index 0) and $9$ (for group B, index 1). We retrieve the group-by keys $[A,B]$ corresponding to the inverse indices $[0,1]$.
@@ 8.1. Get Join Index
Note that if either join column uses RLE, the matching index yields a one-to-many join and needs to be duplicated by the run length. If both join columns use RLE, then each pair of matching runs yields a many-to-many join, and the final run lengths are determined by the product of their lengths (by applying Algorithm 2 to duplicate the runs).
@@
Example 9.
Consider generating Join Indices between a Plain DataColumn and an RLE DataColumn as shown in Figure 6. The Plain DataColumn contains values [A, B, B] at positions 0, 1, and 2. The RLE DataColumn contains values $v=[A,B]$ with start positions $s=[0,2]$, end positions $e=[1,2]$, and run lengths $l=[2,1]$ ($l=e-s+1$).

### v2

@@ 5. Logical Operations
Table 3. Output MaskColumn encoding for AN. The table is symmetric and we only show the upper triangular matrix.
@@ 5.1. AND
  • Plain mask AND Plain mask: Use PyTorch’s logical AND operator (&). For example, $[T,F,T,F]$ AND $[T,T,F,F]$ yields $[T,F,F,F]$.
@@
  • RLE mask AND Plain mask: Direct AND is inefficient due to searching each ‘True’ value in Plain mask. Instead, convert RLE mask to Index (rle_to_index) or Plain (rle_to_plain) then apply AND following Table 2. The choice depends on the selectivity threshold (total elements / selected elements). If this ratio exceeds the threshold, convert to an Index mask; otherwise, convert to a Plain mask. By default, we choose 20, which was determined through offline profiling on our GPU system. This threshold balances RLE compression benefits against conversion costs—providing good performance when RLE is beneficial while keeping conversion overhead acceptable.
  • RLE mask AND Index mask: Check if Index values fall within RLE ranges using PyTorch’s bucketize. Use idx_in_rle or rle_contain_idx depending on relative mask sizes.
  • Plain mask AND Index mask: Select Index values where Plain positions are True using PyTorch’s subscript operator ($[]$).
  • Index mask AND Index mask: Check if values in one Index tensor appear in the other using bucketize on the larger tensor.
@@
Alternative Design. For RLE mask AND Plain mask, our current design only converts RLE to Index/Plain representations, because it prioritizes simplicity and predictable performance. An alternative design could convert Plain to RLE instead. For instance, when a Plain mask undergoes multiple AND operations, it may become highly selective and potentially benefit from RLE compression. However, such conversion could introduce significant overhead.
[…]
However, we acknowledge that, while for a single operator, converting Plain to RLE is too costly, the RLE representation may benefit downstream operations. To decide optimally though is challenging and potentially requires runtime statistics. We recognize there are interesting query optimization opportunities in this space, but leave these encoding decisions as future work.
Figure 4. Performance comparison of alternative AND designs across Plain compression ratios. Alternative Design Plain $\rightarrow$ RLE has faster AND operation when Plain is highly compressible. However, the large overhead of Plain to RLE conversion negates the gain.
@@
[…]
@@
  • Plain mask OR Plain mask: Use PyTorch’s logical OR operator ($|$). For example, $[T,F,T,F]$ OR $[T,T,F,F]$ yields $[T,T,T,F]$.
  • RLE mask OR RLE mask: Compute union of two sorted range lists by identifying start/end points of consecutive True values and determining non-overlapping segments.
  • RLE mask OR Plain mask: Convert the RLE mask to Index (rle_to_index) or Plain (rle_to_plain) then apply OR operation. Similar to AND, the choice depends on the selectivity threshold with default value 20.
  • RLE Mask OR Index Mask: Bucketize based on relative mask sizes using idx_in_rle or rle_contain_idx.
  • Plain Mask OR Index Mask: Use PyTorch’s subscript ([]) operator.
  • Index Mask OR Index Mask: Merge sorted Index tensors using merge_sorted_idx. Bucketize the larger tensor for efficiency, track element origins with flags, then merge using conditional selection. For small tensors, concatenate and sort.
@@ 5.3. NOT
The NOT operator takes a single MaskColumn as input and produces an output MaskColumn.
  • NOT Plain mask: Use PyTorch’s complement ($\sim$) operator for bitwise negation.
  • NOT RLE mask: Use complement_rle primitive to compute gaps between consecutive runs. Each gap starts at previous run end + 1 and ends at next run start - 1. Requires metadata to track total column rows.
  • NOT Index mask: Use complement_index primitive to compute gaps between indices, tracking column size. Output is in RLE format (not Index) because Index masks are sparse, making NOT results continuous and RLE-suited.
Implementation details and examples are provided in Section A.1.
@@ 5.4. Operating on Composite MaskColumns
The previous sections discussed logical operations for non-composite MaskColumns. We now extend these operations to Composite masks (Plain + Index, or RLE + Index). Rather than implementing specialized operators for each composite configuration, we leverage a key insight: Composite MaskColumns can be conceptualized as the disjunction (OR operation) of their constituent mask tensors. This observation enables us to decompose logical operations on Composite masks into sequences of operations on their individual components, thereby reusing our existing non-composite operators. We apply Boolean algebra identities—specifically De Morgan’s Laws (De Morgan, 1847), along with Associative and Distributive properties—to systematically transform these operations. Let $m^{rle}$ and $m^{idx}$ denote the RLE and Index components of Composite mask $m$, respectively.
  • NOT: $\neg(m_{1}^{rle}\vee m_{1}^{idx})=(\neg m_{1}^{rle})\wedge(\neg m_{1}^{idx})$
  • OR: $({\color[rgb]{0.3,0.3,1}m_{1}^{rle}}\vee{\color[rgb]{0.3,0.3,1}m_{1}^{idx}})\vee({\color[rgb]{1,0.3,0.3}m_{2}^{rle}}\vee{\color[rgb]{1,0.3,0.3}m_{2}^{idx}})=({\color[rgb]{0.3,0.3,1}m_{1}^{rle}}\vee{\color[rgb]{1,0.3,0.3}m_{2}^{rle}})\vee({\color[rgb]{0.3,0.3,1}m_{1}^{idx}}\vee{\color[rgb]{1,0.3,0.3}m_{2}^{idx}})$
[…]
    $({\color[rgb]{0.3,0.3,1}m_{1}^{rle}}\wedge{\color[rgb]{1,0.3,0.3}m_{2}^{rle}})\vee({\color[rgb]{0.3,0.3,1}m_{1}^{rle}}\wedge{\color[rgb]{1,0.3,0.3}m_{2}^{idx}})\vee({\color[rgb]{0.3,0.3,1}m_{1}^{idx}}\wedge{\color[rgb]{1,0.3,0.3}m_{2}^{rle}})\vee({\color[rgb]{0.3,0.3,1}m_{1}^{idx}}\wedge{\color[rgb]{1,0.3,0.3}m_{2}^{idx}})$
@@ 6. Arithmetic, Comparison, and Selection Operations
We focus on operators that perform point-wise/position-wise operations on values from two input columns, for positions that are common to both columns. This includes the logical AND operator (Section 5.1), binary arithmetic operators, comparison operators, and selection operators with filter predicates. Joins need additional handling (Section 8). The logical OR operator (Section 5.2) does not fit this category in the general case since the result covers a union of positions from the input columns which can be larger than the set of positions for any one column in the presence of gaps.
The main challenge in performing these operations on compressed inputs is that, unlike Plain encoding, there is no alignment between positional representations of the inputs. Thus, point-wise operations are not directly possible. We need to first align the positional representations of the inputs, including the value tensors as needed, then perform the operations on the aligned segments. We call this Alignment.
@@
Consider $c_{1}+c_{2}$ on input columns $c_{1}$ and $c_{2}$ both in RLE format. $c_{1}^{val}=[4,1,3],c_{1}^{start}=[0,10,20],c_{1}^{end}=[9,19,39]$; $c_{2}^{val}=[6,8],c_{2}^{start}=[0,15],c_{2}^{end}=[14,39]$. Thus, $c_{1}$ has 3 runs (lengths 10, 10, 20) and $c_{2}$ has 2 runs (lengths 15, 25). Similar examples apply for ,,/,<,>,=-,*,/,<,>,= operators.
Due to run misalignment, we cannot directly do $c_{1}^{val}+c_{2}^{val}$. Misalignment can happen even with the same number of runs if run lengths differ. We first align the runs of the input columns, similar to the AND operator (Section 5.1) with an additional step of reconstructing the value tensors. After alignment, we get two RLE columns $r_{1}$ and $r_{2}$ with $r_{1}^{start}=r_{2}^{start}=[0,10,15,20]$, $r_{1}^{end}=r_{2}^{end}=[9,14,19,39]$, $r_{1}^{val}=[4,1,1,3]$, $r_{2}^{val}=[6,6,8,8]$. With identical position tensors, we can do point-wise addition: $r_{1}^{val}+r_{2}^{val}=[10,7,9,11]$. Operations with scalar operands (e.g., $c*2,1-c,c\geq 3$) are simple—no alignment needed, just operate on value tensors for RLE and Index encodings.
For selection operations (e.g., SELECT C FROM D WHERE P), we compute MaskColumn $m$ from predicate P, align $m$ and C, then apply $m$ to the tensor representations of $C$ if needed. For RLE and Index encodings, alignment performs selection (no final application needed). For Plain encodings, final mask application is required.
@@ 7. Group-By-Aggregation Operations
For aggregation queries, the process is decomposed into two phases:
@@
Both phases are conceptually straightforward in PyTorch—function torch.unique can handle grouping by building inverse indices based on unique values in the group-by columns, and function torch.scatter can handle aggregation based on these inverse indices. However, the challenge lies in managing heterogeneous compression schemes across DataColumns. For example, consider a query that groups by two columns, one RLE-compressed and one Plain. The torch.unique function will not work because it requires aligned Plain DataColumns, but the RLE column values are not aligned with the Plain column. This heterogeneity prevents direct application of torch.unique and torch.scatter due to misalignment. We solve this by applying our Alignment technique (Section 6) to both phases.
@@ 7.1. Grouping Phase
The grouping phase constructs an inverse index that maps each row to its corresponding group; rows in the same group share the same data values across the group-by columns. It takes a set of aligned DataColumns for the group-by as input and returns an inverse index. The inverse index is a single tensor containing values from the range 0 to N-1, where N is the number of unique values across the group-by columns. The inverse index length is the same as the number of values in the group-by columns—for Plain data, it is the number of rows; for RLE and Index data, it is the number of RLE runs or index points.
@@ 7.2. Aggregating Phase
To implement aggregation, we apply the torch.scatter function to the data columns, scattering based on the inverse indices. This approach efficiently computes aggregates like SUM, COUNT, AVG, STD, VAR, MIN, and MAX for each group. Note that if the data is RLE-compressed, each value must be repeated based on its corresponding RLE range size. We need to account for this when scattering the data. This depends on the aggregation function. A detailed walkthrough is provided in Section A.2.
@@
  • SUM, COUNT: For RLE columns, instead of expanding the data and counting/summing individual values, we compute the run lengths: l=es+1l=e-s+1. This is sufficient for COUNT. For SUM, we multiply the run lengths by the values (vlv*l).
@@ 8.1. Get Join Index
Note that if either join column uses RLE, the matching index yields a one-to-many join and needs to be duplicated by the run length. If both join columns use RLE, then each pair of matching runs yields a many-to-many join, and the final run lengths are determined by the product of their lengths (by applying Algorithm 2 to duplicate the runs). A detailed example of this join index generation process is provided in Section A.3.